OpenStudy (anonymous):

PLEASE HELP MEDAL WILL BE REWARDED!!!!!! Write the sum using summation notation, assuming the suggested pattern continues. 2 - 6 + 18 - 54 + ... Very Confused...

terenzreignz (terenzreignz):

Look for a pattern... 2 -6 18 -54 This is a hunch... but don't you just keep multiplying a certain number to it? ;)

mathslover (mathslover):

I think they are being multiplied by 3 : So we can write them as : \(x- 3x + 9x - 27x ...\)

OpenStudy (anonymous):

@terenzreignz Yes! -3! so the answer has to be in summation notation form now where do i go ?

OpenStudy (anonymous):

\[\sum_{i=1}^{?}-3(i-1)\]

OpenStudy (anonymous):

it looks wrong though

OpenStudy (anonymous):

@Chelsea04 So close but its not one of the answers! I will give you the 4 answers i can choose from

OpenStudy (anonymous):

the format's probably not right but the idea is there, if you times 2 but -3 you get -6 if you times -6 by -3 you get 18 if you times 18 by -3 you get -54

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}2(-3)^n\]

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}2(3)^n\]

OpenStudy (anonymous):

the first one, i think

OpenStudy (anonymous):

\[\sum_{n=0}^{?\infty}2(3)^(n+1)\]

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}2(-3)^(n-1)\]

OpenStudy (anonymous):

it seems like the first one you wrote

OpenStudy (anonymous):

I think its the last one byt im not sure...

OpenStudy (anonymous):

Ill go with the first one then!! You are probably right! :)

OpenStudy (anonymous):

if you sub in 0 into the last one, the first value you get is 2/-3 or something like that